OpenStudy (anonymous):

how to integrate sqrt(x)/(1+x)?

OpenStudy (anonymous):

try \(u=\sqrt{x}\)

OpenStudy (agent0smith):

\[\Large \int\limits_{}^{}\frac{ \sqrt x }{ 1+x } dx\] Looks like you'll have to make a trig substitution as well as a regular substitution...idk exactly how it'll work out since i haven't tried it yet. Maybe start with u = sqrtx, so u^2 = x, and du= dx/(2sqrtx) = dx/2u so dx = 2u du \[\Large \int\limits_{}^{}\frac{ u }{ 1+u^2 }*2u du\] idk if this is going anywhere... but you can use int. by parts from here i think.

OpenStudy (anonymous):

then \(u^2=x, 2udu=dx\) and integrate \[\int \frac{u^2}{1+u^2}du\]

OpenStudy (anonymous):

oops i dropped a 2

OpenStudy (anonymous):

\[\Large 2\int\frac{ u^2 }{ 1+u^2 } du\]

OpenStudy (dan815):

doh why did i delete that

OpenStudy (anonymous):

\(\frac{u^2}{1+u^2}=\frac{u^2+1-1}{u^2+1}=1-\frac{1}{1+u^2}\)

OpenStudy (anonymous):

now it should be easier

OpenStudy (anonymous):

Satellite73 how did you get that?

OpenStudy (anonymous):

don't need parts, although i guess it could work \[\int \frac{du}{1+u^2}\] is well known

OpenStudy (anonymous):

how did i get what?

OpenStudy (isaiah.feynman):

@satellite73 Good algebra usage there.

OpenStudy (anonymous):

lol thanks you can also divide i guess, but it is easier to do it that way i think

OpenStudy (agent0smith):

you can also get this by doing polynomial long division, btw @ryanmk54 \[\Large \frac{u^2}{1+u^2}=\frac{u^2+1-1}{u^2+1}=1-\frac{1}{1+u^2}\]

OpenStudy (isaiah.feynman):

Wow..the algebra simplifies everything easily.

OpenStudy (anonymous):

\[u^2 /(1+u^2)\] I got u over\[u/(u^2 +1)\]

OpenStudy (anonymous):

don't forget that \(2udu=dx\) which is why you have \(u^2\) up top and not \(u\)

OpenStudy (anonymous):

Did you then simplify by creatively adding 0? \[u^2 + 1 -1\]

OpenStudy (agent0smith):

^yes he did!

OpenStudy (anonymous):

yes, but that is because i have seen this gimmick many times without it, you have to divide since you have \(u^2\) up top and \(u^2\) in the bottom, you know it will be \(1+\frac{a}{1+u^2}\) so it is easier just to force it instead of diving using long division

OpenStudy (anonymous):

*dividing

OpenStudy (anonymous):

Thx agent0smith and satellite73

OpenStudy (anonymous):

Integral of the second term would just be ln(1-u^2)du?

OpenStudy (agent0smith):

^no, integral of 1/(1+x^2) = arctan x ...iirc

OpenStudy (anonymous):

Your right thx